# Architecting Scalable Trapped Ion Quantum Computers using Surface Codes

### v1

@@ 5.1. Performance and Noise Models
Determining the logical error rate of the code is more involved and requires a noise simulation. We use Stim simulations for this purpose (Gidney, 2021). The input to Stim is a hardware noise model, which in our case is a realistic error-model for QCCD systems based on modelling of the relevant noise sources in trapped-ion hardware, as described in (Gutiérrez et al., 2019). In addition, the model has been modified to account for the dependence of qubit gate error rates on the vibrational energy of ions, the number of ions, and the gate duration, as outlined in (Murali et al., 2020).
@@
    where $t$ is the duration of the operation and $T_{2}=2.2$ seconds is the coherence time for the trapped-ion qubit, obtained from real experiments that demonstrated its accuracy (Gutiérrez et al., 2019).
@@
Errors from ion movement are incorporated into the fidelity model, obtained from (Murali et al., 2020), which influences the probabilities of errors $e_{2}$ and $e_{3}$. The fidelity of the qubit gate is influenced by two primary factors: background heating from the trap’s electromagnetic field and thermal motion from higher vibrational energy of the ion chain. The fidelity $p(e_{2}),p(e_{3})$ is expressed as:
@@
We have validated our parameters against hardware data sheets from Quantinuum and IonQ. We also consider a range of gate improvements (1X to 10X) in our experiments to account for future improvements. A 5X improvement in our setup corresponds to $\approx 10^{-3}$ depolarising error rates per qubit gate, which is comparable to the best-known devices from Quantinuum and IonQ (Moses et al., 2023a; Chen et al., 2024).
Cooling Model: Cooling ions before qubit gates decrease physical error rates at the expense of increased execution times. To model the effect of cooling in the WISE wiring method, the noise model in Table 1 is modified, setting the baseline two-qubit gate error to $2\times 10^{-3}$ and the one-qubit gate error to $3\times 10^{-3}$, while ignoring heating effects by adding an extra 850 $\mu s$ to the two-qubit gate time (Pino et al., 2021).
@@ 5.2. Resource Estimation Model
  • The number of dynamic/shim electrodes per zone: $N_{lz/de}=10$, $N_{jz/de}=20$, and $N_{se/z}=10$ (Malinowski et al., 2023).
Decreasing the trap capacity increases the number of electrodes for a fixed qubit count. This is because the ratio of junction zones to linear zones, $N_{jz}/N_{lz}$ increases, so lower trap capacities require more electrodes (since junction zones require more electrodes than linear zones) (Malinowski et al., 2023).
@@
In the WISE architecture, the data rate is $\approx 50\,\text{Mbit/s}\times N_{\text{DACs}}$, where the number of DACs is $N_{\text{DACs}}\approx 100+\frac{N_{se}}{100}$, while $N_{se}$ denotes the number of shim electrodes. As a result, the WISE architecture scales two orders of magnitude more favourably in terms of data rate compared to the standard architecture, significantly reducing the burden on control electronics (Malinowski et al., 2023).
@@ 6.2. Architecture configurations
We explore trap capacities, ranging from 2 to 30, along with the grid, switch and linear connectivities. We also explore the standard choice for control system wiring, where each DAC is connected to one electrode, and the WISE architecture (Malinowski et al., 2023). Since our study aims to understand the design choices for future systems with potentially improved hardware, we scale the physical error rates by a factor called gate improvement. For example, a 10X gate improvement corresponds to every gate having a 10X lower physical error rate and the dephasing physical error rate on idling qubits being 10X less. The gate improvement in our experiment varies from 1X to 10X.
@@ 6.3. Metrics
Logical Error Rate: The logical error rate quantifies the primary objective of QEC: suppressing quantum errors to levels that enable fault-tolerant computation. The experiment looks to identify configurations capable of achieving a $10^{-9}$ logical error rate, which is a minimum requirement for large-scale algorithms (Acharya et al., 2023).
@@ 6.4. Logical Error Rate Calculation Using Stim
The logical error rate calculation is performed by interfacing the physical noise model and the execution schedule of the compiled circuit into a noisy quantum circuit in Stim (Gidney, 2021). We use Stim version 1.13.0.
@@ 6.5. Baselines
Our QEC compiler (implemented in Python 3.11) is benchmarked against two other trapped-ion QCCD compilers: QCCDSim (Murali et al., 2020) and Muzzle The Shuttle (Saki et al., 2022) in terms of ion movement time and number of movement operations.
@@ 7.1. Accurate and Scalable QEC Compiler
Table 3 compares the performance our QEC compiler QCCDSim (Murali et al., 2020) and Muzzle The Shuttle (Saki et al., 2022). We benchmarked five rounds of error correction to account for any changes in qubit layout across rounds. For all baselines, the time required to execute gates is the same. Therefore, we focus on movement time (time required for ion reconfigurations) and the number of movement operations. Our QEC compiler achieves an average 3.85X reduction in movement time and an average 1.91X reduction in movement operations compared to the best of the two compilers in each test case. In the best case, the improvement is up to 6.03X. For the 2D rotated surface code, the QEC compiler successfully compiles five rounds of error correction across a wide range of trap capacities and code distances. In contrast, QCCDSim and MuzzleTheShuttle either produce suboptimal schedules or fail to compile entirely, especially at higher code distances. These results show that our compiler is well-suited for architectural evaluations.
@@ 7.3. Choice of Trap Capacity
Unlike prior NISQ studies, which recommend the use of traps with capacity in the range of 20-30 ions (Murali et al., 2020), we advocate the use of a trap capacity of two to obtain logical qubits with hardware efficiency, low error rates, and a constant runtime regardless of code distance.
@@ 7.4. Choice of wiring method
A key power bottleneck in the standard architecture is that each electrode is wired to a separate DAC. WISE (Malinowski et al., 2023) overcomes this with a more intelligent wiring mechanism, trading off execution time for reduced power consumption. Which mechanism is the most suitable for logical qubit implementation? Figure 13(a) compares the data required for WISE and the standard wiring mechanism. For the standard mechanism, we only use trap capacity 2. Whereas, for WISE, we examine trap capacities ranging from 2 to 30 but only show the curves for three capacities, since the trends are similar at other capacities. Compared to the standard architecture, WISE achieves an improvement of more than two orders of magnitude in data rate (and, therefore, in power consumption).
@@
Figure 13(b) compares the elapsed time at different logical error rates. For the WISE architecture, the elapsed time scales in proportion to the desired logical error rate. For every 10X improvement desired in the logical error rate, the elapsed time increases by 1.17X. WISE suffers from limited transport flexibility, allowing only one transport operation at a time. Under an odd-even sort global reconfiguration scheme (Malinowski et al., 2023), this limitation results in logical clock speeds that are up to 2525X longer than those of standard QPUs, for logical error rates near $10^{-9}$. This runtime increase is acceptable for near-term fault-tolerant applications such as quantum dynamics (van Dam et al., 2024). However, for large applications such as factoring, which already require month-long computations on trapped ion systems(Lekitsch et al., 2017), such a runtime increase will lead to impractical executions that run over a year.
@@ 8. Related Work
This work builds on previous advancements in QCCD system architecture and QEC optimisation. For instance, Gutiérrez et al. (Gutiérrez et al., 2019) inspire the test infrastructure to validate executable QCCD circuits. The relevance of compiler-driven architectural co-design for QCCD systems is demonstrated by Murali et al. (Murali et al., 2020), which examines the influence of micro-architectural choices on the performance of NISQ algorithms. Similarly, Wu et al. (Wu et al., 2022) address the challenges in bridging quantum hardware and QEC codes by proposing a framework for efficient implementation and optimisation of surface codes for superconducting architectures. This study extends these concepts by tailoring a QEC compiler to the specific demands of QCCD-based systems, aiming to provide a systematic approach to co-designing hardware and software for fault-tolerant quantum computing.
While there exist QCCD compilers for QEC other than the two benchmarked in (§7.1), such as the MQTIonShuttler (Schoenberger et al., 2024), we do not benchmark our compiler against these, since they assume distinct memory and processing zones in their QCCD architecture, which is not suitable for surface code implementation. TISCC (Leblond et al., 2023) fixes the trap capacity as two and the standard grid topology (Lekitsch et al., 2017), then compiles and simulates high-level logical circuits into a quantum circuit on physical qubits using the surface code. The compiler does not map to primitive QCCD directly but uses the performance models of these primitives for resource estimation.
Consideration of Limiting Factors: In contrast to superconducting platforms, decoder runtimes are not the limiting factor for ion-trap systems since their cycle time is considerably longer. Specialist hardware is already available for the fast decoding of surface codes up to a distance of 8 (Barber et al., 2025).
@@ 9. Conclusion
TI qubit technology is at the threshold of supporting systems with several logical qubits. Current demonstrations of logical qubits are limited to small systems of less than 60 physical qubits. To scale up to systems with several hundred physical qubits (tens of logical qubits), we need to understand what the right trap capacities and topologies are and how control systems must be designed to support QEC workloads. The TI community has been exploring these choices for several years, with 1) monolithic, large trap capacity devices (e.g., IonQ Forte) 2) QCCD devices with small trap capacities (e.g., Quantinuum H2) 3) architecture research showing the value of QCCD systems with 15-25 ions per trap (Murali et al., 2020) and 4) other manual design efforts (Malinowski et al., 2023; Lekitsch et al., 2017; Valentini et al., 2024).

### v2

@@ 5.1. Performance and Noise Models
Determining the logical error rate of the code is more involved and requires a noise simulation. We use Stim simulations for this purpose (gidney2021stim). The input to Stim is a hardware noise model, which in our case is a realistic error-model for QCCD systems based on modelling of the relevant noise sources in trapped-ion hardware, as described in (transversality_lattice_surgery). In addition, the model has been modified to account for the dependence of qubit gate error rates on the vibrational energy of ions, the number of ions, and the gate duration, as outlined in (murali2020architectingnoisyintermediatescaletrapped).
@@
    where $t$ is the duration of the operation and $T_{2}=2.2$ seconds is the coherence time for the trapped-ion qubit, obtained from real experiments that demonstrated its accuracy (transversality_lattice_surgery).
@@
Errors from ion movement are incorporated into the fidelity model, obtained from (murali2020architectingnoisyintermediatescaletrapped), which influences the probabilities of errors $e_{2}$ and $e_{3}$. The fidelity of the qubit gate is influenced by two primary factors: background heating from the trap’s electromagnetic field and thermal motion from higher vibrational energy of the ion chain. The fidelity $p(e_{2}),p(e_{3})$ is expressed as:
@@
We have validated our parameters against hardware data sheets from Quantinuum and IonQ. We also consider a range of gate improvements (1X to 10X) in our experiments to account for future improvements. A 5X improvement in our setup corresponds to $\approx 10^{-3}$ depolarising error rates per qubit gate, which is comparable to the best-known devices from Quantinuum and IonQ (H2Racetrack; Chen_2024).
Cooling Model: Cooling ions before qubit gates decrease physical error rates at the expense of increased execution times. To model the effect of cooling in the WISE wiring method, the noise model in Table 1 is modified, setting the baseline two-qubit gate error to $2\times 10^{-3}$ and the one-qubit gate error to $3\times 10^{-3}$, while ignoring heating effects by adding an extra 850 $\mu s$ to the two-qubit gate time (Pino_2021).
@@ 5.2. Resource Estimation Model
  • The number of dynamic/shim electrodes per zone: $N_{lz/de}=10$, $N_{jz/de}=20$, and $N_{se/z}=10$ (1000qubits).
Decreasing the trap capacity increases the number of electrodes for a fixed qubit count. This is because the ratio of junction zones to linear zones, $N_{jz}/N_{lz}$ increases, so lower trap capacities require more electrodes (since junction zones require more electrodes than linear zones) (1000qubits).
@@
In the WISE architecture, the data rate is $\approx 50\,\text{Mbit/s}\times N_{\text{DACs}}$, where the number of DACs is $N_{\text{DACs}}\approx 100+\frac{N_{se}}{100}$, while $N_{se}$ denotes the number of shim electrodes. As a result, the WISE architecture scales two orders of magnitude more favourably in terms of data rate compared to the standard architecture, significantly reducing the burden on control electronics (1000qubits).
@@ 6.2. Architecture configurations
We explore trap capacities, ranging from 2 to 30, along with the grid, switch and linear connectivities. We also explore the standard choice for control system wiring, where each DAC is connected to one electrode, and the WISE architecture (1000qubits). Since our study aims to understand the design choices for future systems with potentially improved hardware, we scale the physical error rates by a factor called gate improvement. For example, a 10X gate improvement corresponds to every gate having a 10X lower physical error rate and the dephasing physical error rate on idling qubits being 10X less. The gate improvement in our experiment varies from 1X to 10X.
@@ 6.3. Metrics
Logical Error Rate: The logical error rate quantifies the primary objective of QEC: suppressing quantum errors to levels that enable fault-tolerant computation. The experiment looks to identify configurations capable of achieving a $10^{-9}$ logical error rate, which is a minimum requirement for large-scale algorithms (suppressing_quantum_errors_by_scaling_surface_code).
@@ 6.4. Logical Error Rate Calculation Using Stim
The logical error rate calculation is performed by interfacing the physical noise model and the execution schedule of the compiled circuit into a noisy quantum circuit in Stim (gidney2021stim). We use Stim version 1.13.0.
@@ 6.5. Baselines
Our QEC compiler (implemented in Python 3.11) is benchmarked against two other trapped-ion QCCD compilers: QCCDSim (murali2020architectingnoisyintermediatescaletrapped) and Muzzle The Shuttle (muzzletheshuttle) in terms of ion movement time and number of movement operations.
@@ 7.1. Accurate and Scalable QEC Compiler
Table 3 compares the performance our QEC compiler QCCDSim (murali2020architectingnoisyintermediatescaletrapped) and Muzzle The Shuttle (muzzletheshuttle). We benchmarked five rounds of error correction to account for any changes in qubit layout across rounds. For all baselines, the time required to execute gates is the same. Therefore, we focus on movement time (time required for ion reconfigurations) and the number of movement operations. Our QEC compiler achieves an average 3.85X reduction in movement time and an average 1.91X reduction in movement operations compared to the best of the two compilers in each test case. In the best case, the improvement is up to 6.03X. For the 2D rotated surface code, the QEC compiler successfully compiles five rounds of error correction across a wide range of trap capacities and code distances. In contrast, QCCDSim and MuzzleTheShuttle either produce suboptimal schedules or fail to compile entirely, especially at higher code distances. These results show that our compiler is well-suited for architectural evaluations.
@@ 7.3. Choice of Trap Capacity
Unlike prior NISQ studies, which recommend the use of traps with capacity in the range of 20-30 ions (murali2020architectingnoisyintermediatescaletrapped), we advocate the use of a trap capacity of two to obtain logical qubits with hardware efficiency, low error rates, and a constant runtime regardless of code distance.
@@ 7.4. Choice of wiring method
A key power bottleneck in the standard architecture is that each electrode is wired to a separate DAC. WISE (1000qubits) overcomes this with a more intelligent wiring mechanism, trading off execution time for reduced power consumption. Which mechanism is the most suitable for logical qubit implementation? Figure 13(a) compares the data required for WISE and the standard wiring mechanism. For the standard mechanism, we only use trap capacity 2. Whereas, for WISE, we examine trap capacities ranging from 2 to 30 but only show the curves for three capacities, since the trends are similar at other capacities. Compared to the standard architecture, WISE achieves an improvement of more than two orders of magnitude in data rate (and, therefore, in power consumption).
@@
Figure 13(b) compares the elapsed time at different logical error rates. For the WISE architecture, the elapsed time scales in proportion to the desired logical error rate. For every 10X improvement desired in the logical error rate, the elapsed time increases by 1.17X. WISE suffers from limited transport flexibility, allowing only one transport operation at a time. Under an odd-even sort global reconfiguration scheme (1000qubits), this limitation results in logical clock speeds that are up to 2525X longer than those of standard QPUs, for logical error rates near $10^{-9}$. This runtime increase is acceptable for near-term fault-tolerant applications such as quantum dynamics (vandam2024usingazurequantumresource). However, for large applications such as factoring, which already require month-long computations on trapped ion systems(Lekitsch_2017), such a runtime increase will lead to impractical executions that run over a year.
@@ 8. Related Work
This work builds on previous advancements in QCCD system architecture and QEC optimisation. For instance, Gutiérrez et al. (transversality_lattice_surgery) inspire the test infrastructure to validate executable QCCD circuits. The relevance of compiler-driven architectural co-design for QCCD systems is demonstrated by Murali et al. (murali2020architectingnoisyintermediatescaletrapped), which examines the influence of micro-architectural choices on the performance of NISQ algorithms. Similarly, Wu et al. (synthesis_framework_stitching) address the challenges in bridging quantum hardware and QEC codes by proposing a framework for efficient implementation and optimisation of surface codes for superconducting architectures. This study extends these concepts by tailoring a QEC compiler to the specific demands of QCCD-based systems, aiming to provide a systematic approach to co-designing hardware and software for fault-tolerant quantum computing.
While there exist QCCD compilers for QEC other than the two benchmarked in (§7.1), such as the MQTIonShuttler (schoenberger2024shuttlingscalabletrappedionquantum), we do not benchmark our compiler against these, since they assume distinct memory and processing zones in their QCCD architecture, which is not suitable for surface code implementation. TISCC (Leblond_2023) fixes the trap capacity as two and the standard grid topology (Lekitsch_2017), then compiles and simulates high-level logical circuits into a quantum circuit on physical qubits using the surface code. The compiler does not map to primitive QCCD directly but uses the performance models of these primitives for resource estimation.
Consideration of Limiting Factors: In contrast to superconducting platforms, decoder runtimes are not the limiting factor for ion-trap systems since their cycle time is considerably longer. Specialist hardware is already available for the fast decoding of surface codes up to a distance of 8 (Barber_2025).
@@ 9. Conclusion
TI qubit technology is at the threshold of supporting systems with several logical qubits. Current demonstrations of logical qubits are limited to small systems of less than 60 physical qubits. To scale up to systems with several hundred physical qubits (tens of logical qubits), we need to understand what the right trap capacities and topologies are and how control systems must be designed to support QEC workloads. The TI community has been exploring these choices for several years, with 1) monolithic, large trap capacity devices (e.g., IonQ Forte) 2) QCCD devices with small trap capacities (e.g., Quantinuum H2) 3) architecture research showing the value of QCCD systems with 15-25 ions per trap (murali2020architectingnoisyintermediatescaletrapped) and 4) other manual design efforts (1000qubits; Lekitsch_2017; valentini2024demonstrationtwodimensionalconnectivityscalable).